OpenStudy (anonymous):

What is a polynomial in standard form with zeros at 4,-1,1

OpenStudy (anonymous):

polynomial would be in the form: y = (x - 4)(x - 1)(x + 1) = (x - 4)(x^2 - 1) = y = x^3 - 4x^2 - x + 4